# A dynamic circuit for the honeycomb Floquet code

### v3

@@ III A review of the honeycomb Floquet code
A subtlety arises when considering logical operators in the Floquet code. For any logical degree of freedom, no fixed representative of the logical operator commutes with all gauge measurements. Instead, we modify the logical representative at each time step, multiplying it by gauge operators that have just been measured to form a new representative that commutes with the next gauge measurement. We illustrate this process for a horizontal/vertical logical operator pair in 2. While the measurement cycle has period $3$, the logical representatives have period $6$.
@@ IV Standard vs dynamic circuit for the Floquet code
Figure 3: Top: Standard circuit for performing XXXX gauge measurements, using one ancilla per gauge operator. The circuits for YYYY and $ZZ$ are similar. Bottom: Gauge measurements in the dynamic circuit. We measure $XX$ gauge operators by shrinking them to a single qubit, which we then measure. In the YYYY and ZZZZ gauge measurements we alternate the measured qubits to remove leakage. In the next three gauge measurements (not shown), we reverse the target and control of each gate and continue the alternating measurement pattern. The two circuits have the same number of two-qubit gates, resets, and measurements.
We present two circuits for measuring the gauge operators, the standard ancilla-based circuit and our new dynamic one, illustrated in Fig. 3. The standard circuit requires one qubit per gauge operator, while the dynamic one requires no ancilla qubits and operates by shrinking the gauge operator onto a single qubit and measuring that qubit directly. Note that in the surface code, switching to the dynamic circuit meant we could only measure half as many operators at once (but got twice as much information per measurement). In contrast, in the Floquet code we get improved measurements while measuring all $XX$ operators simultaneously, which is the source of improved performance.
@@
Our dynamic circuit thus requires fewer measurement cycles per logical operation. However, every layer of our dynamic circuit must be executed sequentially, meaning a full cycle takes time $(6t_{\text{gate}}+3t_{\text{meas}}+3t_{\text{reset}})$. This contrasts with the standard circuit, in which the first layer of gates to measure $YY$ commutes with the last layer of gates to measure $XX$, and measured qubits do not need to be immediately reset. The standard circuit can therefore be compressed; in [gidney2021fault] they provide a circuit that takes amortized time $(3t_{\text{gate}}+3\max[t_{\text{meas}},t_{\text{reset}}])$ provided the hardware can execute measure and reset operations in parallel. As a result, depending on hardware constraints, a logical operation in the dynamic circuit may still be slower.
@@ Appendix A Noise Model
  • Single-qubit gates are followed by a single-qubit depolarizing channel with strength $p$.
  • Two-qubit gates are followed by a two-qubit depolarizing channel with strength $p$.
  • Idle qubits during single or two-qubit gates incur a single-qubit depolarizing channel with strength $p$.
@@
  • Measurement results are flipped with probability $p$.
@@ Appendix B Timelike distance
We have numerically verified this behavior using STIM [gidney2021stim]. For fixed $d$, we generate circuits with a varying number $N_{m}$ of measurement rounds and produce their detector error models. We can then numerically find the minimum distance between detectors occuring around the beginning of the circuit (some fixed distance $\delta N_{m}$ from 0) and detectors occuring around the end of the circuit (some fixed distance $\delta N_{m}$ from $N_{m}$). The precise value of $N_{m}$ is not particularly important, as long as we keep it fixed as we scale $N_{m}$. We can then consider two quantities: (1) $d_{\text{graph}}$, the minimum number of graphlike errors it takes to connect the beginning detectors to end detectors (2) $d_{\text{hyper}}$, the minimum number of graphlike errors connecting the detectors if we decompose all hyperedges $h$ into complete graphs on the degree-$|h|$ vertices of $h$. These distances provide upper/lower bounds on the timelike distance $d_{t}$: $d_{\text{graph}}\geq d_{t}\geq d_{\text{hyper}}$. For the standard circuit we find $d_{\text{graph}}=d_{\text{hyper}}=3N_{m}/4$, so that we require $N_{m}=4d_{t}/3$ to reach distance $d_{t}$. For the dynamic circuit, we find $d_{\text{graph}}=2N_{m},d_{\text{hyper}}=N_{m}$, so we require at most $N_{m}=d_{t}$ to reach distance $d_{t}$ (but $N_{m}=d_{t}/2$ may be sufficient). Establishing the precise relationship between $d_{t}$ and $N_{m}$ for the dynamic circuit will require a stability simulation [gidney2022stability].
@@ Appendix C Additional numerical results
Here, we include a plot of our extrapolation of logical error rate as a function of distance, used to compute the teraquop footprint for $p=10^{-3}$ as well as graphs of the separate horizontal and vertical logical error rates, rather than summed as in the main text. We present results from both MWPM decoding, as well as the more accurate but slower correlated matching decoding [fowler2013optimal]. The V logical error probability is the probability that the vertical logical operator is incorrect, i.e. the probability that an undetected horizontal logical operator has been applied, and similarly for the H logical error probability. The larger V logical error probability for the standard circuit was also seen in [gidney2022benchmarking]. The steeper slope of the V logical error probability in the dynamic circuit results from a complicated issue of error multiplicity that is difficult to fully account for. We note that all of these probabilities have a slope steeper than $p^{\lfloor d/2\rfloor}$ despite the circuits having a numerically-verified graphlike distance of $d$, due to higher-distance errors dominating at these values of $p$.

### v4

@@ III A review of the honeycomb Floquet code
A subtlety arises when considering logical operators in the Floquet code. For any logical degree of freedom, no fixed representative of the logical operator commutes with all gauge measurements. Instead, we modify the logical representative at each time step, multiplying it by gauge operators that have just been measured to form a new representative that commutes with the next gauge measurement. We illustrate this process for a horizontal/vertical logical operator pair in 2. While the measurement cycle has period 33, the logical representatives have period 66.
@@ IV Standard vs dynamic circuit for the Floquet code
Figure 3: Top: Standard circuit for performing XXXX gauge measurements, using one ancilla per gauge operator. The circuits for YYYY and ZZZZ are similar. Bottom: Gauge measurements in the dynamic circuit. We measure XXXX gauge operators by shrinking them to a single qubit, which we then measure. In the YYYY and ZZZZ gauge measurements we alternate the measured qubits to remove leakage. In the next three gauge measurements (not shown), we reverse the target and control of each gate and continue the alternating measurement pattern. The two circuits have the same number of two-qubit gates, resets, and measurements.
We present two circuits for measuring the gauge operators, the standard ancilla-based circuit and our new dynamic one, illustrated in Fig. 3. The standard circuit requires one qubit per gauge operator, while the dynamic one requires no ancilla qubits and operates by shrinking the gauge operator onto a single qubit and measuring that qubit directly. Note that in the surface code, switching to the dynamic circuit meant we could only measure half as many operators at once (but got twice as much information per measurement). In contrast, in the Floquet code we get improved measurements while measuring all XXXX operators simultaneously, which is the source of improved performance.
@@
Our dynamic circuit thus requires fewer measurement cycles per logical operation. However, every layer of our dynamic circuit must be executed sequentially, meaning a full cycle takes time $(6t_{\text{gate}}+3t_{\text{meas}}+3t_{\text{reset}})$. This contrasts with the standard circuit, in which the first layer of gates to measure YYYY commutes with the last layer of gates to measure XXXX, and measured qubits do not need to be immediately reset. The standard circuit can therefore be compressed; in [gidney2021fault] they provide a circuit that takes amortized time $(3t_{\text{gate}}+3\max[t_{\text{meas}},t_{\text{reset}}])$ provided the hardware can execute measure and reset operations in parallel. As a result, depending on hardware constraints, a logical operation in the dynamic circuit may still be slower.
@@ Appendix A Noise Model
  • Single-qubit gates are followed by a single-qubit depolarizing channel with strength pp.
  • Two-qubit gates are followed by a two-qubit depolarizing channel with strength pp.
  • Idle qubits during single or two-qubit gates incur a single-qubit depolarizing channel with strength pp.
@@
  • Measurement results are flipped with probability pp.
@@ Appendix B Timelike distance
We have numerically verified this behavior using STIM [gidney2021stim]. For fixed $d$, we generate circuits with a varying number $N_{m}$ of measurement rounds and produce their detector error models. We can then numerically find the minimum distance between detectors occuring around the beginning of the circuit (some fixed distance $\delta N_{m}$ from 0) and detectors occuring around the end of the circuit (some fixed distance $\delta N_{m}$ from $N_{m}$). The precise value of $\delta N_{m}$ is not particularly important, as long as we keep it fixed as we scale $N_{m}$. We can then consider two quantities: (1) $d_{\text{graph}}$, the minimum number of graphlike errors it takes to connect the beginning detectors to end detectors (2) $d_{\text{hyper}}$, the minimum number of graphlike errors connecting the detectors if we decompose all hyperedges $h$ into complete graphs on the degree-$|h|$ vertices of $h$. These distances provide upper/lower bounds on the timelike distance $d_{t}$: $d_{\text{graph}}\geq d_{t}\geq d_{\text{hyper}}$. For the standard circuit we find $d_{\text{graph}}=d_{\text{hyper}}=3N_{m}/4$, so that we require $N_{m}=4d_{t}/3$ to reach distance $d_{t}$. For the dynamic circuit, we find $d_{\text{graph}}=2N_{m},d_{\text{hyper}}=N_{m}$, so we require at most $N_{m}=d_{t}$ to reach distance $d_{t}$ (but $N_{m}=d_{t}/2$ may be sufficient). Establishing the precise relationship between $d_{t}$ and $N_{m}$ for the dynamic circuit will require a stability simulation [gidney2022stability].
@@ Appendix C Additional numerical results
Here, we include a plot of our extrapolation of logical error rate as a function of distance, used to compute the teraquop footprint for $p=10^{-3}$ as well as graphs of the separate horizontal and vertical logical error rates, rather than summed as in the main text. We present results from both MWPM decoding, as well as the more accurate but slower correlated matching decoding [fowler2013optimal]. The V logical error probability is the probability that the vertical logical operator is incorrect, i.e. the probability that an undetected horizontal logical operator has been applied, and similarly for the H logical error probability. The larger V logical error probability for the standard circuit was also seen in [gidney2022benchmarking]. The steeper slope of the V logical error probability in the dynamic circuit results from a complicated issue of error multiplicity that is difficult to fully account for. We note that all of these probabilities have a slope steeper than $p^{\lfloor d/2\rfloor}$ despite the circuits having a numerically-verified graphlike distance of $d$, due to higher-distance errors dominating at these values of pp.